3/4x + 2/3y=7
3/5x - 1/2y=18
solve by elimination or substitution
I have done this problem 15 times and cannot get the answer. I am doing something wrong and need to see the steps involved more than just the right answer because I don't know what I have done wrong.

Answer by scott8148(6628)

multiply 1st eqn by 60 (3 * 4 * 5) ___ 45x + 40y = 420 (A)

multiply 2nd eqn by 80 ___ 48x - 40y = 1440 (B)

add eqn (A) and eqn (B), eliminating y ___ 93x = 1860

find x and substitute back to find y